How many solutions in non-negative integers are there to the equation x1 + x2+ x3 + x4 = 19


1
Expert's answer
2021-11-05T14:41:12-0400

Theorem

Let "n" and "k" are positive integers. Then the number of non-negative integer solutions of the equation "x_1+x_2+...+x_n=k" is given by "C(k+n-1, k)."

Given "n=4, k=19."

Then the number of non-negative integer solutions of the equation

"x_1+x_2+x_3+x_4=19"


is


"C(19+4-1, 19)=\dbinom{22}{19}=\dfrac{22!}{3!(22-19)!}"

"=\dfrac{22(21)(20)}{1(2)(3)}=1540"

There are 1540 non-negative integer solutions of the equation "x_1+x_2+x_3+x_4=19."